# Evolution of Discordance

### v1

@@ 3.1 Attempts at capturing co-evolution
What happens when the evolutions of the voter model and the graph are locked in a mutual feed-back? Early work for dense graphs, based on simulations and intuitive reasoning, was carried out by Holme and Newman [16] and by Durrett, Gleeson, Lloyd and Varghese [12]. Depending on the choice of dynamics, a phase transition may occur between convergence to a final state where only one opinion survives (= consensus) or both opinions survive (= polarisation).
@@
Basu and Sly [7] showed that in Models (I) and (II) with high probability, i.e., with probability tending to $1$ as $n\to\infty$, the following dichotomy hold: for $\beta$ small, $\tau$ is of order $n^{2}$ and polarisation occurs at equal fractions of the two opinions, while for $\beta$ large, $\tau$ is of order $n^{3}$ and polarisation occurs at different fractions of the two opinions. As $\beta\to\infty$ the fraction of the minority opinion tends to zero, i.e., there is asymptotic consensus. Moreover, in Model (I) for fixed $\beta$ with high probability the fraction of the minority opinion at time $\tau$ is bounded away from zero. Durrett, Gleeson, Lloyd and Varghese [12] anticipated these results with the help of simulations and approximate computations, and predicted a sharp threshold between the two types of behaviour.
@@
Holme and Newman [16] conjectured a similar phase transition as that conjectured in Durrett, Gleeson, Lloyd and Varghese [12], based on simulation. Note that in this model the dynamics also ends when there are no edges between vertices with different opinions.

### v2

@@ 3.1 Attempts at capturing co-evolution
What happens when the evolutions of the voter model and the graph are locked in a mutual feed-back? Early work for dense graphs, based on simulations and intuitive reasoning, was carried out by Holme and Newman [16] and by Durrett, Gleeson, Lloyd, Mucha, Shi, Sivakoff, Socolar and Varghese [12]. Depending on the choice of dynamics, a phase transition may occur between convergence to a final state where only one opinion survives (= consensus) or both opinions survive (= polarisation).
@@
Basu and Sly [7] showed that in Models (I) and (II) with high probability, i.e., with probability tending to $1$ as $n\to\infty$, the following dichotomy hold: for $\beta$ small, $\tau$ is of order $n^{2}$ and polarisation occurs at equal fractions of the two opinions, while for $\beta$ large, $\tau$ is of order $n^{3}$ and polarisation occurs at different fractions of the two opinions. As $\beta\to\infty$ the fraction of the minority opinion tends to zero, i.e., there is asymptotic consensus. Moreover, in Model (I) for fixed $\beta$ with high probability the fraction of the minority opinion at time $\tau$ is bounded away from zero. These results were anticipated in [12] with the help of simulations and approximate computations, and led to the prediction of a sharp threshold between the two types of behaviour.
@@
Holme and Newman [16] conjectured a similar phase transition as that conjectured in [12], based on simulation. Note that in this model the dynamics also ends when there are no edges between vertices with different opinions.